# Learning Theory for Kernel Bilevel Optimization

### v1

@@ Appendix D Generalization Properties
D.1 Point-wise estimates
We present a point-wise upper-bound on the value error ${\left\lvert\mathcal{F}(\omega)-\widehat{\mathcal{F}}(\omega)\right\rvert}$ and gradient error ${\left\lVert\nabla\mathcal{F}(\omega)-\widehat{\nabla\mathcal{F}}(\omega)\right\rVert}$. To this end, we introduce the following notation for the error between the inner and outer objectives and their empirical approximations evaluated at the optimal inner solution $h_{\omega}^{\star}$:
@@ D.1 Point-wise estimates
By abuse of notation, we introduce the following error between partial derivatives of $L_{in}$ and $\widehat{L}_{in}$ (resp. $L_{out}$ and $\widehat{L}_{out}$), evaluated at $(\omega,h_{\omega}^{\star})$, i.e.,
$\partial_{h}\delta_{\omega}^{out}$ $\coloneqq{\left\lVert\partial_{h}L_{out}(\omega,h^{\star}_{\omega})-\partial_{h}\widehat{L}_{out}(\omega,h^{\star}_{\omega})\right\rVert}_{\mathcal{H}},\quad$ $\partial_{\omega}\delta_{\omega}^{out}$ $\coloneqq{\left\lVert\partial_{\omega}L_{out}(\omega,h^{\star}_{\omega})-\partial_{\omega}\widehat{L}_{out}(\omega,h^{\star}_{\omega})\right\rVert},$
$\partial_{h}\delta_{\omega}^{in}$ $\coloneqq{\left\lVert\partial_{h}L_{in}(\omega,h^{\star}_{\omega})-\partial_{h}\widehat{L}_{in}(\omega,h^{\star}_{\omega})\right\rVert}_{\mathcal{H}},\quad$ $\partial_{\omega}\delta_{\omega}^{in}$ $\coloneqq{\left\lVert\partial_{\omega}L_{in}(\omega,h^{\star}_{\omega})-\partial_{\omega}\widehat{L}_{in}(\omega,h^{\star}_{\omega})\right\rVert},$
$\partial_{h}^{2}\delta_{\omega}^{in}$ $\coloneqq{\left\lVert\partial_{h}^{2}L_{in}(\omega,h^{\star}_{\omega})-\partial_{h}^{2}\widehat{L}_{in}(\omega,h^{\star}_{\omega})\right\rVert}_{\operatorname{op}},\quad$ $\partial_{\omega,h}^{2}\delta_{\omega}^{in}$ $\coloneqq{\left\lVert\partial_{\omega,h}^{2}L_{in}(\omega,h^{\star}_{\omega})-\partial_{\omega,h}^{2}\widehat{L}_{in}(\omega,h^{\star}_{\omega})\right\rVert}_{\operatorname{op}}.$
Proposition D.1.
Under Assumptions (B), (D), (A) and (C), the following holds for any $\omega\in\Omega$:
@@
Let $\omega\in\Omega$. The function $h\mapsto\widehat{L}_{in}(\omega,h)$ is $\lambda$-strongly convex and Fréchet differentiable by Propositions A.3 and A.1. Moreover, $\hat{h}_{\omega}$ is the minimizer of $h\mapsto\widehat{L}_{in}(\omega,h)$ by definition. Therefore, using Lemma G.2, we obtain a control on the distance in $\mathcal{H}$ to the optimum $\hat{h}_{\omega}$ of $h\mapsto\widehat{L}_{in}(\omega,h)$ in terms of the gradient $\partial_{h}\widehat{L}_{in}(\omega,h)$:
@@
In particular, choosing $h=h^{\star}_{\omega}$ yields the first inequality. The fact that ${\left\lVert\partial_{h}\widehat{L}_{in}(\omega,h_{\omega}^{\star})\right\rVert}_{\mathcal{H}}=\partial_{h}\delta_{\omega}^{in}$ follows from the optimality of $h_{\omega}^{\star}$ which implies that $\partial_{h}L_{in}(\omega,h_{\omega}^{\star})=0$. ∎
Proposition D.2.
Under Assumptions (B), (D), (A) and (C), the following inequalities hold for any $\omega\in\Omega$:
@@
where $M_{out}$, $\bar{M}_{out}$, and $\bar{M}_{in}$ are the positive constants defined in Propositions C.3 and C.4.
@@
Lipschitz continuity of some functions of interest. Let $\omega\in\Omega$. According to Proposition C.1, both $h^{\star}_{\omega}(x)$ and $\hat{h}_{\omega}(x)$ lie in the compact interval $\mathcal{V}\coloneqq\left[-\frac{B\kappa}{\lambda},\frac{B\kappa}{\lambda}\right]\subset\mathbb{R}$, for any $x\in\mathcal{X}$, where $B\coloneqq\sup_{\omega\in\text{hull}(\Omega),y\in\mathcal{Y}}\left|\partial_{v}\ell_{in}(\omega,0,y)\right|>0$. By Assumption (B), $\mathcal{Y}$ is a compact set. Hence $\Omega\times\mathcal{V}\times\mathcal{Y}$ is a compact set as well. Furthermore, by Assumption (C), $(\omega,v,y)\mapsto\ell_{in}(\omega,v,y)$, $(\omega,v,y)\mapsto\ell_{out}(\omega,v,y)$, and their derivatives are all continuous over the compact domain $\Omega\times\mathcal{V}\times\mathcal{Y}$. Therefore, these functions and their derivatives are bounded on this domain. In particular, this also holds when $v$ takes the specific values $h^{\star}_{\omega}(x)$ or $\hat{h}_{\omega}(x)$. Let $\bar{v}$ be either $h^{\star}_{\omega}(x)$ or $\hat{h}_{\omega}(x)$, for any $x\in\mathcal{X}$. For any $\omega\in\Omega$, and $y\in\mathcal{Y}$, we have:
@@
${\left\lVert\partial_{v}\partial_{\omega,v}^{2}\ell_{in}(\omega,\bar{v},y)\right\rVert}_{\operatorname{op}}$ $\leq\sup_{\omega\in\Omega,v\in\mathcal{V},y\in\mathcal{Y}}{\left\lVert\partial_{\omega}\partial_{v}^{2}\ell_{in}(\omega,v,y)\right\rVert}\leq\bar{M}_{in}<+\infty$
@@
Upper-bounds. We have:
@@
where the first line uses the definition of $(\omega,h)\mapsto\widehat{L}_{out}(\omega,h)$, the second line applies the triangle inequality, the third line leverages the fact that $v\mapsto\ell_{out}(\omega,v,y)$ is $M_{out}$-Lipschitz continuous, for any $\omega\in\Omega$ and $y\in\mathcal{Y}$, and the last line follows from the reproducing property of the RKHS $\mathcal{H}$, Cauchy-Schwarz’s inequality, and Assumption (A) to bound ${\left\lVert K(x,\cdot)\right\rVert}_{\mathcal{H}}$ by $\sqrt{\kappa}$. Similarly, we obtain:
$\partial_{h}E_{\omega}^{out}\leq\bar{M}_{out}\kappa{\left\lVert h^{\star}_{\omega}-\hat{h}_{\omega}\right\rVert}_{\mathcal{H}},\quad\partial_{\omega}E_{\omega}^{out}\leq\bar{M}_{out}\sqrt{\kappa}{\left\lVert h^{\star}_{\omega}-\hat{h}_{\omega}\right\rVert}_{\mathcal{H}},\quad\partial_{h}^{2}E_{\omega}^{in}\leq\bar{M}_{in}\kappa\sqrt{\kappa}{\left\lVert h^{\star}_{\omega}-\hat{h}_{\omega}\right\rVert}_{\mathcal{H}},$
$\partial_{\omega,h}^{2}E_{\omega}^{in}\leq\bar{M}_{in}\kappa{\left\lVert h^{\star}_{\omega}-\hat{h}_{\omega}\right\rVert}_{\mathcal{H}}.$
@@
Proposition D.3.
Under Assumptions (B), (D), (A) and (C), the following inequalities hold for any $\omega\in\Omega$:
@@
where $C_{out}$ and $C_{in}$ are the positive constants defined in Proposition D.2.
@@
Upper-bound on ${\left\lVert\partial_{h}L_{out}(\omega,h_{\omega}^{\star})\right\rVert}_{\mathcal{H}}$. We have:
@@
where the first line follows from Proposition A.1, the second line results from the triangle inequality, and the last line uses Assumption (A) to bound ${\left\lVert K(x,\cdot)\right\rVert}_{\mathcal{H}}$ by $\sqrt{\kappa}$. Furthermore, we know by Proposition C.1 that $(\omega,h_{\omega}^{\star}(x),y)$ belongs to the compact subset $\Omega\times\mathcal{V}\times\mathcal{Y}$ and by Proposition C.3 that $\partial_{v}\ell_{out}(\omega,h^{\star}_{\omega}(x),y)$ is bounded by a constant $M_{out}$ on $\text{hull}(\Omega)\times\mathcal{V}\times\mathcal{Y}$. Hence, it follows that:
@@
where $C_{out}$ is defined in Proposition D.2.
Upper-boud on $\quad{\left\lVert\partial^{2}_{\omega,h}L_{in}(\omega,h_{\omega}^{\star})\right\rVert}_{\operatorname{op}}$. According to Proposition A.2, $\partial_{\omega,h}^{2}L_{in}(\omega,h^{\star}_{\omega})$ is a Hilbert-Schmidt operator, which points to:
$\left\|\partial_{\omega,h}^{2}L_{in}(\omega,h^{\star}_{\omega})\right\|_{\operatorname{op}}\leq\left\|\partial_{\omega,h}^{2}L_{in}(\omega,h^{\star}_{\omega})\right\|_{\operatorname{HS}}=\sqrt{\sum_{l=1}^{d}\left\|\partial_{\omega_{k},h}^{2}L_{in}(\omega,h^{\star}_{\omega})\right\|_{\mathcal{H}}^{2}}.$ (15)
This means that to find an upper-bound on $\left\|\partial_{\omega,h}^{2}L_{in}(\omega,h^{\star}_{\omega})\right\|_{\operatorname{op}}$, it suffices to establish an upper-bound on $\left\|\partial_{\omega_{l},h}^{2}L_{in}(\omega,h^{\star}_{\omega})\right\|_{\mathcal{H}}^{2}$ for any $l\in\{1,\ldots,d\}$. For a fixed $l\in\{1,\ldots,d\}$, we have:
@@
where the first line follows from Proposition A.2, the second line is a consequence of Jensen’s inequality applied on the convex function $\|\cdot\|^{2}$, and the last line applies Assumption (A) to bound ${\left\lVert K(x,\cdot)\right\rVert}_{\mathcal{H}}^{2}$ by $\kappa$. Incorporating this upper-bound into Equation 15 yields:
$\left\|\partial_{\omega,h}^{2}L_{in}(\omega,h^{\star}_{\omega})\right\|_{\operatorname{op}}\leq\sqrt{\mathbb{E}_{\mathbb{P}}\left[\left\|\partial_{\omega,v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)\right\|^{2}\right]d\kappa}\leq M_{in}\sqrt{d\kappa}\leq C_{out},$
where we used Proposition C.3 to bound $\partial_{\omega,v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)$ by the constant $M_{in}$.
Upper-bound on ${\left\lVert\partial^{2}_{\omega,h}\widehat{L}_{in}(\omega,\hat{h}_{\omega})\right\rVert}_{\operatorname{op}}$. The derivation of this upper bound follows the same steps as the previous one, with the only differences being the use of $\widehat{L}_{in}$ instead of $L_{in}$, and $\hat{h}_{\omega}$ instead of $h^{\star}_{\omega}$.
Note that in the last step of each of the three upper-bounds, we used the fact that the functions we are dealing with are continuous (by Assumption (C)) on $\Omega\times\mathcal{V}\times\mathcal{Y}$, which is compact because $\Omega$ and $\mathcal{Y}$ are compact by Assumption (B) and $\mathcal{V}$ is a compact interval of $\mathbb{R}$ defined in Proposition C.1. Hence, those functions are bounded. ∎
Proposition D.4 (Approximation bounds).
Under Assumptions (B), (D), (A) and (C), the following holds for any $\omega\in\Omega$:
@@
${\left\lVert\nabla\mathcal{F}(\omega)-\widehat{\nabla\mathcal{F}}(\omega)\right\rVert}\leq\partial_{\omega}\delta_{\omega}^{out}+\frac{C_{in}}{\lambda}\partial_{h}\delta_{\omega}^{out}+\frac{C_{out}C_{in}}{\lambda^{2}}\partial_{h}^{2}\delta_{\omega}^{in}+\frac{C_{out}}{\lambda}\partial_{\omega,h}^{2}\delta_{\omega}^{in}+\frac{C_{out}}{\lambda}{\left(1+2\frac{C_{in}}{\lambda}+\frac{C_{in}^{2}}{\lambda^{2}}\right)}\partial_{h}\delta_{\omega}^{in},$
where the constants $C_{in}$ and $C_{out}$ are given in Proposition D.2.
@@
${\left\lvert\mathcal{F}(\omega)-\hat{\mathcal{F}}(\omega)\right\rvert}\leq\underbrace{{\left\lvert L_{out}(\omega,h^{\star}_{\omega})-\widehat{L}_{out}(\omega,h^{\star}_{\omega})\right\rvert}}_{\delta_{\omega}^{out}}+\underbrace{{\left\lvert\widehat{L}_{out}(\omega,h^{\star}_{\omega})-\widehat{L}_{out}(\omega,\hat{h}_{\omega})\right\rvert}}_{E_{\omega}^{out}},$ (16)
According to Proposition D.2, the error term $E_{\omega}^{out}$ is controlled by the norm of the difference $h^{\star}_{\omega}-\hat{h}_{\omega}$, i.e., $E_{\omega}^{out}\leq C_{out}{\left\lVert h^{\star}_{\omega}-\hat{h}_{\omega}\right\rVert}_{\mathcal{H}}$. Moreover, by Proposition D.1, we know that ${\left\lVert h^{\star}_{\omega}-\hat{h}_{\omega}\right\rVert}_{\mathcal{H}}\leq\frac{1}{\lambda}\partial_{h}\delta_{\omega}^{in}$. Therefore, combining both bounds yields: $E_{\omega}^{out}\leq\frac{C_{out}}{\lambda}\partial_{h}\delta_{\omega}^{in}$. The upper-bound on value function follows by substituting the previous inequality into Equation 16.
Control on the gradient. By Proposition A.4, we have the following expression for the total gradient $\nabla\mathcal{F}$:
@@
^$\widehat{\nabla\mathcal{F}}(\omega)=\partial_{\omega}\widehat{L}_{out}(\omega,\hat{h}_{\omega})-\partial_{\omega,h}^{2}\widehat{L}_{in}(\omega,\hat{h}_{\omega})\left(\partial_{h}^{2}\widehat{L}_{in}(\omega,\hat{h}_{\omega})\right)^{-1}\partial_{h}\widehat{L}_{out}(\omega,\hat{h}_{\omega}).$
To simplify notations, for any $h\in\mathcal{H}$, we introduce the following operators $R(h),\hat{R}(h):\mathcal{H}\to\Omega$:
$R(h)=\partial_{\omega,h}^{2}L_{in}(\omega,h)\left(\partial_{h}^{2}L_{in}(\omega,h)\right)^{-1}\quad\text{and}\quad\hat{R}(h)=\partial_{\omega,h}^{2}\widehat{L}_{in}(\omega,h)\left(\partial_{h}^{2}\widehat{L}_{in}(\omega,h)\right)^{-1}.$
@@
$\nabla\mathcal{F}(\omega)-\widehat{\nabla\mathcal{F}}(\omega)=$ $\left(\partial_{\omega}L_{out}(\omega,h^{\star}_{\omega})-\partial_{\omega}\widehat{L}_{out}(\omega,h^{\star}_{\omega})\right)+\left(\partial_{\omega}\widehat{L}_{out}(\omega,h^{\star}_{\omega})-\partial_{\omega}\widehat{L}_{out}(\omega,\hat{h}_{\omega})\right)$
$-\hat{R}(\hat{h}_{\omega}){\left(\left(\partial_{h}L_{out}(\omega,h^{\star}_{\omega})-\partial_{h}\widehat{L}_{out}(\omega,h^{\star}_{\omega})\right)+\left(\partial_{h}\widehat{L}_{out}(\omega,h^{\star}_{\omega})-\partial_{h}\widehat{L}_{out}(\omega,\hat{h}_{\omega})\right)\right)}$
$-\left(R(h^{\star}_{\omega})-\hat{R}(\hat{h}_{\omega})\right)\partial_{h}L_{out}(\omega,h^{\star}_{\omega}).$
@@
By taking the norm of the above equality and using triangle inequality, we obtain the following upper-bound:

### v3

@@ Appendix E Generalization Properties
E.1 Point-wise estimates
We present a point-wise upper bound on the value error $\left\lvert\mathcal{F}(\omega)-\widehat{\mathcal{F}}(\omega)\right\rvert$ and gradient error $\left\lVert\nabla\mathcal{F}(\omega)-\widehat{\nabla\mathcal{F}}(\omega)\right\rVert$. To this end, we introduce the following notation for the error between the inner and outer objectives and their empirical approximations evaluated at the optimal inner solution $h_{\omega}^{\star}$:
@@ E.1 Point-wise estimates
By abuse of notation, we introduce the following errors between partial derivatives of $L_{in}$ and $\widehat{L}_{in}$ (resp. $L_{out}$ and $\widehat{L}_{out}$), evaluated at $(\omega,h_{\omega}^{\star})$, i.e.,
$\partial_{h}\delta_{\omega}^{out}\coloneqq\left\lVert\partial_{h}L_{out}(\omega,h^{\star}_{\omega})-\partial_{h}\widehat{L}_{out}(\omega,h^{\star}_{\omega})\right\rVert_{\mathcal{H}},\partial_{\omega}\delta_{\omega}^{out}\coloneqq\left\lVert\partial_{\omega}L_{out}(\omega,h^{\star}_{\omega})-\partial_{\omega}\widehat{L}_{out}(\omega,h^{\star}_{\omega})\right\rVert,$
$\partial_{h}\delta_{\omega}^{in}\coloneqq\left\lVert\partial_{h}L_{in}(\omega,h^{\star}_{\omega})-\partial_{h}\widehat{L}_{in}(\omega,h^{\star}_{\omega})\right\rVert_{\mathcal{H}},\partial_{\omega}\delta_{\omega}^{in}\coloneqq\left\lVert\partial_{\omega}L_{in}(\omega,h^{\star}_{\omega})-\partial_{\omega}\widehat{L}_{in}(\omega,h^{\star}_{\omega})\right\rVert,$
$\partial_{h}^{2}\delta_{\omega}^{in}\coloneqq\left\lVert\partial_{h}^{2}L_{in}(\omega,h^{\star}_{\omega})-\partial_{h}^{2}\widehat{L}_{in}(\omega,h^{\star}_{\omega})\right\rVert_{\operatorname{op}},\partial_{\omega,h}^{2}\delta_{\omega}^{in}\coloneqq\left\lVert\partial_{\omega,h}^{2}L_{in}(\omega,h^{\star}_{\omega})-\partial_{\omega,h}^{2}\widehat{L}_{in}(\omega,h^{\star}_{\omega})\right\rVert_{\operatorname{op}}.$
Proposition E.1.
Under Assumptions˜(A), (C), (E), (B) and (D), the following holds for any $\omega\in\Omega$:
@@
Let $\omega\in\Omega$. The function $h\mapsto\widehat{L}_{in}(\omega,h)$ is $\lambda$-strongly convex and Fréchet differentiable by Propositions˜B.3 and B.1. Moreover, $\hat{h}_{\omega}$ is the minimizer of $h\mapsto\widehat{L}_{in}(\omega,h)$ by definition. Therefore, using Lemma˜H.2, we obtain a control on the distance in $\mathcal{H}$ to the optimum $\hat{h}_{\omega}$ of $h\mapsto\widehat{L}_{in}(\omega,h)$ in terms of the gradient $\partial_{h}\widehat{L}_{in}(\omega,h)$:
@@
In particular, choosing $h=h^{\star}_{\omega}$ yields the inequality. The fact that $\left\lVert\partial_{h}\widehat{L}_{in}(\omega,h_{\omega}^{\star})\right\rVert_{\mathcal{H}}=\partial_{h}\delta_{\omega}^{in}$ follows from the optimality of $h_{\omega}^{\star}$ which implies that $\partial_{h}L_{in}(\omega,h_{\omega}^{\star})=0$. ∎
Proposition E.2.
Under Assumptions˜(A), (C), (E), (B) and (D), the following inequalities hold for any $\omega\in\Omega$:
@@
where $M_{out}$, $\bar{M}_{out}$, and $\bar{M}_{in}$ are the positive constants defined in Propositions˜D.3 and D.4.
@@
Lipschitz continuity of some functions of interest. Let $\omega\in\Omega$. According to Proposition˜D.1, both $h^{\star}_{\omega}(x)$ and $\hat{h}_{\omega}(x)$ lie in the compact interval $\mathcal{V}\coloneqq\left[-\frac{B\kappa}{\lambda},\frac{B\kappa}{\lambda}\right]\subset\mathbb{R}$, for any $x\in\mathcal{X}$, where $B\coloneqq\sup_{\omega\in\text{hull}(\Omega),y\in\mathcal{Y}}\left|\partial_{v}\ell_{in}(\omega,0,y)\right|>0$. By Assumption˜(C), $\mathcal{Y}$ is a compact set. Hence, $\Omega\times\mathcal{V}\times\mathcal{Y}$ is a compact set as well. Furthermore, by Assumption˜(D), $(\omega,v,y)\mapsto\ell_{in}(\omega,v,y)$, $(\omega,v,y)\mapsto\ell_{out}(\omega,v,y)$, and their derivatives are all continuous over the compact domain $\Omega\times\mathcal{V}\times\mathcal{Y}$. Therefore, these functions and their derivatives are bounded on this domain. In particular, this also holds when $v$ takes the specific values $h^{\star}_{\omega}(x)$ or $\hat{h}_{\omega}(x)$. Let $\bar{v}$ be either $h^{\star}_{\omega}(x)$ or $\hat{h}_{\omega}(x)$, for any $x\in\mathcal{X}$. For any $\omega\in\Omega$ and $y\in\mathcal{Y}$, we have:
@@
$\left\lVert\partial_{v}\partial_{\omega,v}^{2}\ell_{in}(\omega,\bar{v},y)\right\rVert_{\operatorname{op}}$ $\leq\sup_{\omega\in\Omega,v\in\mathcal{V},y\in\mathcal{Y}}\left\lVert\partial_{\omega}\partial_{v}^{2}\ell_{in}(\omega,v,y)\right\rVert\leq\bar{M}_{in}<+\infty.$
@@
Upper bounds. We have:
@@
where the first line uses the definition of $(\omega,h)\mapsto\widehat{L}_{out}(\omega,h)$, the second line applies the triangle inequality, the third line leverages the fact that $v\mapsto\ell_{out}(\omega,v,y)$ is $M_{out}$-Lipschitz continuous, for any $\omega\in\Omega$ and $y\in\mathcal{Y}$, and the last line follows from the reproducing property of the RKHS $\mathcal{H}$, Cauchy-Schwarz’s inequality, and Assumption˜(B) to bound $\left\lVert K(x,\cdot)\right\rVert_{\mathcal{H}}$ by $\sqrt{\kappa}$. Similarly, we obtain:
$\partial_{h}E_{\omega}^{out}\leq\bar{M}_{out}\kappa\left\lVert h^{\star}_{\omega}-\hat{h}_{\omega}\right\rVert_{\mathcal{H}},\quad\partial_{\omega}E_{\omega}^{out}\leq\bar{M}_{out}\sqrt{\kappa}\left\lVert h^{\star}_{\omega}-\hat{h}_{\omega}\right\rVert_{\mathcal{H}},$
$\partial_{h}^{2}E_{\omega}^{in}\leq\bar{M}_{in}\kappa\sqrt{\kappa}\left\lVert h^{\star}_{\omega}-\hat{h}_{\omega}\right\rVert_{\mathcal{H}},\quad\partial_{\omega,h}^{2}E_{\omega}^{in}\leq\bar{M}_{in}\kappa\left\lVert h^{\star}_{\omega}-\hat{h}_{\omega}\right\rVert_{\mathcal{H}}.$
@@
Proposition E.3.
Under Assumptions˜(A), (C), (E), (B) and (D), the following inequalities hold for any $\omega\in\Omega$:
@@
where $C_{out}$ and $C_{in}$ are the positive constants defined in Proposition˜E.2.
@@
Upper bound on $\left\lVert\partial_{h}L_{out}(\omega,h_{\omega}^{\star})\right\rVert_{\mathcal{H}}$. We have:
@@
where the first line follows from Proposition˜B.1, the second line results from the triangle inequality, and the last line uses Assumption˜(B) to bound $\left\lVert K(x,\cdot)\right\rVert_{\mathcal{H}}$ by $\sqrt{\kappa}$. Furthermore, we know by Proposition˜D.1 that $(\omega,h_{\omega}^{\star}(x),y)$ belongs to the compact subset $\Omega\times\mathcal{V}\times\mathcal{Y}$, and by Proposition˜D.3 that $\partial_{v}\ell_{out}(\omega,h^{\star}_{\omega}(x),y)$ is bounded by a constant $M_{out}$ on $\text{hull}(\Omega)\times\mathcal{V}\times\mathcal{Y}$. Hence, it follows that:
@@
where $C_{out}$ is defined in Proposition˜E.2.
Upper bound on $\left\lVert\partial^{2}_{\omega,h}L_{in}(\omega,h_{\omega}^{\star})\right\rVert_{\operatorname{op}}$. According to Proposition˜B.2, $\partial_{\omega,h}^{2}L_{in}(\omega,h^{\star}_{\omega})$ is a Hilbert-Schmidt operator, which points to:
$\left\|\partial_{\omega,h}^{2}L_{in}(\omega,h^{\star}_{\omega})\right\|_{\operatorname{op}}\leq\left\|\partial_{\omega,h}^{2}L_{in}(\omega,h^{\star}_{\omega})\right\|_{\operatorname{HS}}=\sqrt{\sum_{l=1}^{d}\left\|\partial_{\omega_{l},h}^{2}L_{in}(\omega,h^{\star}_{\omega})\right\|_{\mathcal{H}}^{2}}.$ (14)
This means that to find an upper bound on $\left\|\partial_{\omega,h}^{2}L_{in}(\omega,h^{\star}_{\omega})\right\|_{\operatorname{op}}$, it suffices to establish an upper bound on $\left\|\partial_{\omega_{l},h}^{2}L_{in}(\omega,h^{\star}_{\omega})\right\|_{\mathcal{H}}^{2}$ for any $l\in\{1,\ldots,d\}$. For a fixed $l\in\{1,\ldots,d\}$, we have:
@@
where the first line follows from Proposition˜B.2, the second line is a consequence of Jensen’s inequality applied on the convex function $\|\cdot\|^{2}$, and the last line applies Assumption˜(B) to bound $\left\lVert K(x,\cdot)\right\rVert_{\mathcal{H}}^{2}$ by $\kappa$. Incorporating this upper bound into Equation˜14 yields:
$\left\|\partial_{\omega,h}^{2}L_{in}(\omega,h^{\star}_{\omega})\right\|_{\operatorname{op}}\leq\sqrt{\mathbb{E}_{\mathbb{P}}\left[\left\|\partial_{\omega,v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)\right\|^{2}\right]d\kappa}\leq M_{in}\sqrt{d\kappa}\leq C_{in},$
where we used Proposition˜D.3 to bound $\partial_{\omega,v}^{2}\ell_{in}(\omega,h^{\star}_{\omega}(x),y)$ by the constant $M_{in}$.
Upper bound on $\left\lVert\partial^{2}_{\omega,h}\widehat{L}_{in}(\omega,\hat{h}_{\omega})\right\rVert_{\operatorname{op}}$. The derivation of this upper bound follows the same steps as the previous one, with the only differences being the use of $\widehat{L}_{in}$ instead of $L_{in}$, and $\hat{h}_{\omega}$ instead of $h^{\star}_{\omega}$.
Note that in the last step of each of the three upper bounds, we used the fact that the functions we are dealing with are continuous by Assumption˜(D) on $\Omega\times\mathcal{V}\times\mathcal{Y}$, which is compact because $\Omega$ is compact, $\mathcal{Y}$ is compact by Assumption˜(C), and $\mathcal{V}$ is a compact interval of $\mathbb{R}$ defined in Proposition˜D.1. Hence, those functions are bounded. ∎
Proposition E.4 (Approximation bounds).
Under Assumptions˜(A), (C), (E), (B) and (D), the following holds for any $\omega\in\Omega$:
@@
$\left\lVert\nabla\mathcal{F}(\omega)-\widehat{\nabla\mathcal{F}}(\omega)\right\rVert\leq$ $\partial_{\omega}\delta_{\omega}^{out}+\frac{C_{in}}{\lambda}\partial_{h}\delta_{\omega}^{out}+\frac{C_{out}C_{in}}{\lambda^{2}}\partial_{h}^{2}\delta_{\omega}^{in}$
$+\frac{C_{out}}{\lambda}\partial_{\omega,h}^{2}\delta_{\omega}^{in}+\frac{C_{out}}{\lambda}\left(1+2\frac{C_{in}}{\lambda}+\frac{C_{in}^{2}}{\lambda^{2}}\right)\partial_{h}\delta_{\omega}^{in},$
where the constants $C_{in}$ and $C_{out}$ are given in Proposition˜E.2.
@@
$\left\lvert\mathcal{F}(\omega)-\widehat{\mathcal{F}}(\omega)\right\rvert\leq\underbrace{\left\lvert L_{out}(\omega,h^{\star}_{\omega})-\widehat{L}_{out}(\omega,h^{\star}_{\omega})\right\rvert}_{\delta_{\omega}^{out}}+\underbrace{\left\lvert\widehat{L}_{out}(\omega,h^{\star}_{\omega})-\widehat{L}_{out}(\omega,\hat{h}_{\omega})\right\rvert}_{E_{\omega}^{out}}.$ (15)
According to Proposition˜E.2, the error term $E_{\omega}^{out}$ is controlled by the norm of the difference $h^{\star}_{\omega}-\hat{h}_{\omega}$, i.e., $E_{\omega}^{out}\leq C_{out}\left\lVert h^{\star}_{\omega}-\hat{h}_{\omega}\right\rVert_{\mathcal{H}}$. Moreover, by Proposition˜E.1, we know that $\left\lVert h^{\star}_{\omega}-\hat{h}_{\omega}\right\rVert_{\mathcal{H}}\leq\frac{1}{\lambda}\partial_{h}\delta_{\omega}^{in}$. Therefore, combining both bounds yields $E_{\omega}^{out}\leq\frac{C_{out}}{\lambda}\partial_{h}\delta_{\omega}^{in}$. The upper bound on the value function follows by substituting the previous inequality into Equation˜15.
Control on the gradient. By Proposition˜B.4, we have the following expression for the total gradient $\nabla\mathcal{F}$:
@@
$\widehat{\nabla\mathcal{F}}(\omega)=\partial_{\omega}\widehat{L}_{out}(\omega,\hat{h}_{\omega})-\partial_{\omega,h}^{2}\widehat{L}_{in}(\omega,\hat{h}_{\omega})\left(\partial_{h}^{2}\widehat{L}_{in}(\omega,\hat{h}_{\omega})\right)^{-1}\partial_{h}\widehat{L}_{out}(\omega,\hat{h}_{\omega}).$
To simplify notations, for any $h\in\mathcal{H}$, we introduce the following operators $R(h),\widehat{R}(h):\mathcal{H}\to\Omega$:
$R(h)=\partial_{\omega,h}^{2}L_{in}(\omega,h)\left(\partial_{h}^{2}L_{in}(\omega,h)\right)^{-1}\quad\text{and}\quad\widehat{R}(h)=\partial_{\omega,h}^{2}\widehat{L}_{in}(\omega,h)\left(\partial_{h}^{2}\widehat{L}_{in}(\omega,h)\right)^{-1}.$
@@
$\nabla\mathcal{F}(\omega)-\widehat{\nabla\mathcal{F}}(\omega)$
$=$ $\left(\partial_{\omega}L_{out}(\omega,h^{\star}_{\omega})-\partial_{\omega}\widehat{L}_{out}(\omega,h^{\star}_{\omega})\right)+\left(\partial_{\omega}\widehat{L}_{out}(\omega,h^{\star}_{\omega})-\partial_{\omega}\widehat{L}_{out}(\omega,\hat{h}_{\omega})\right)$
$-\widehat{R}(\hat{h}_{\omega})\left(\left(\partial_{h}L_{out}(\omega,h^{\star}_{\omega})-\partial_{h}\widehat{L}_{out}(\omega,h^{\star}_{\omega})\right)+\left(\partial_{h}\widehat{L}_{out}(\omega,h^{\star}_{\omega})-\partial_{h}\widehat{L}_{out}(\omega,\hat{h}_{\omega})\right)\right)$
$-\left(R(h^{\star}_{\omega})-\widehat{R}(\hat{h}_{\omega})\right)\partial_{h}L_{out}(\omega,h^{\star}_{\omega}).$
@@
By taking the norm of the above equality and using the triangle inequality, we obtain the following upper bound: